# Iodine and Bromine Radical Reactions in Atmospheric Mercury Oxidation

### v1

@@ IV.2 Atmospheric Mixing Ratios in the Arctic and Antarctic
When looking at the atmospheric concentrations, high maximum mixing ratios of bromine, chlorine, and iodine species have been measured at depletion events both in the Arctic and in the Antarctic. Here maximum concentrations of BrO, IO, and ClO radicals have been measured up to 50 ppt,[85] 30 ppt,[86] and 44 ppt.[87] It is hard to compare these maximum concentrations, since they are all measured from single or rarely occurring episodes. At a few locations, atmospheric concentrations of two or three halogen monoxides have been measured at the same time and place. Saiz-Lopez et al.[88] made measurements of BrO and IO at Hally Bay in the Antarctic and reported similar mixing ratios of BrO and IO with maximums of 20.2 and 20.5 ppt and minimums of both below the detection limit of 2 ppt.
In the dissertation of Zielcke,[31] there are reports of rare events with mixing ratios of BrO over 20 ppt. In the Arctic in 2014, average measurements were made and compared with the solar zenith angle (SZA) with values from 5 ppt (SZA[-85]) in the spring dropping to 1 ppt (SZA[-45]) in the summer and rising again to about 5 ppt in the fall. In the Antarctic, BrO mixing rate was reported over the year with the result of mixing ratios raising from below the detection limit to 3 ppt in the spring (SZA [-90 to -80]), falling again in the summer to 1 ppt and rising to 1.5 ppt in the fall, before dropping to back below the detection limit in the winter.
In the same series of measurements, IO was measured in the Arctic ranging from 0.1 ppt (SZA[90]) spring to 0.3 ppt (SZA[50]) summer. In the Antarctic, the average ranges from 0.35 ppt spring to 0.15 ppt in summer and back to about 0.35 ppt in the fall.[31]
NO in the Arctic dropped from 0.2 ppb in the spring to under the detection limit in the summer, rising again to 0.2 ppb in the fall. In the Antarctic, the mole fraction rises to 1.3 ppb in the spring, dropping below the detection limit in the summer, and rises up back over 1 ppb in the fall.[31]
During the SOLVE[89] science flights over the Norwegian Sea near the Arctic, the atmospheric concentrations of HO, ClO, and NO radicals were measured with average concentrations in clouds of 0.36 ppt, 20.1 ppt, and 24 ppt. In cloud-free surroundings, the measurements were 0.25 ppt, 11.8 ppt, and 18.4 ppt. From fall to winter, HO concentrations fall from about 2 ppt to about 0.1 ppt. ClO radicals falls from about 30 ppt to about 10 ppt. Reasons for the difference in NO mixing ratios compared to Zielcke[31] can be due to local factors such as maritime traffic in the Norwegian Sea.
The average concentrations of OH during the SOLVE mission was 0.026 ppt and was dropping from fall to winter from about 0.1 ppt to about 0.01 ppt.[89]
Few measurements are available concerning the concentrations of the atomic radicals. The only example of the atomic iodine radical I found in literature was measured at Mace Head in Ireland. There measured mixing ratios ranging from under the limit of detection (5.3 ppt) up to 22 ppt were detected on 29 Aug 2007.[90] Atomic chlorine radicals have been estimated to go up to concentrations of about 0.05 ppt.[87] The concentration of atomic bromine radicals measured in Borrow Alaska by Wang et al.[91] ranged from 4 ppt to 14 ppt over an Ozone Depletion Event, with measured Br/BrO ratios ranging from 7.5 during an Ozone Depletion Event to more normal levels of 0.1.
Local conditions vary widely, and the atmospheric concentration of halogen radicals have a significant effect on the AMDE. From Tables 10 and 11 we can only get a partial understanding of these concentrations, making it even harder to predict how external factors such as global warming will change these conditions. There is substantial evidence to suggest that iodine concentrations may rise due to global warming. However, the reactions investigated in this study indicate that the radical species of Br and OBr are the primary contributors. Should iodine species, such as atomic iodine, become more abundant than concentrations of atmospheric bromine species - contrary to current observations - iodine could then compete with bromine in atmospheric reactions during the mercury depletion events. Iodine will also play a role in activating bromine [5]. In this context, it is mainly atomic iodine radicals that will be competitive, whereas OI radicals would need to be substantially more abundant to compete effectively. Based on these observations, we can resolve that the AMDEs are dependent on the concentration of the radical species, and the role of iodine species as contributors therefore cannot be ruled out.
| Species | Maximum | Location | Ref. | |
| mixing ratio ppt | ||||
| BrO | 50 | Polarstern | Antarctic | Wagner 2007 [85] |
| 20 | Alert Base | Arctic | Zielcke 2015 [31] | |
| 20 | Halley Bay | Antarctic | Saiz-Lopez 2015 [41] | |
| IO | 30 | Mace Head | Ireland | Commane 2011 [86] |
| 20 | Halley Bay | Antarctic | Saiz-Lopez 2007 [88] | |
| 3.4 | Hudson’s Bay | Arctic | Mahaja 2010 [92] | |
| 0.8 | Svalbard | Arctic | Wittrock 2000 [93] | |
| ClO | 44 | Barrow | Alaska | Custard 2016 [87] |
| Br | 14 | Barrow | Alaska | Wang 2019 [91] |
| I | 22 | Mace Head | Ireland | Bale 2008 [90] |
| Species | Average yearly variations ppt | location | region | reff. |
| BrO | 5 - 2 | Alert Base | Arctic | Zielcke 2015 [31] |
| 3 - 1 | Scott Base | Antarctic | Zielcke 2015 [31] | |
| IO | 0.3 - 0.1 | Alert Base | Arctic | Zielcke 2015 [31] |
| 0.35 - 0.15 | Scott Base | Antarctic | Zielcke 2015 [31] | |
| ClO | 30 - 10 | SOLVE | Arctic | Simpas 2012 [89] |
| HO | 2 - 0.1 | SOLVE | Arctic | Simpas 2012 [89] |
| NO | 0.0002 - LOD | Alert Base | Arctic | Zielcke 2015 [31] |
| 0.0013 - LOD | Scott Base | Antarctic | Zielcke 2015 [31] | |
| OH | 0.1 - 0.01 | SOLVE | Arctic | Simpas 2012 [89] |
@@ V Conclusion and Outlook
The following observations emerge for the reaction energy. The reactions with the largest stabilization of Gibbs free energy creating Hg(II) from reaction R2 (XHg +Y XHgY) are the reaction forming HgBr with a stabilization energy of 61.09 kcal/mol.
The stabilization by the Gibbs free energy decreases with increasing atomic number, from XHgOI over XHgOBr to XHgOCl. Furthermore, at the other end of the series of molecules, IHgOX molecules are consistently less stable than their BrHgOX counterparts. The least stable XHgOX-product is IHgOCl with a stabilization of 36.06 kcal/mol in Gibbs free energy.
Lower stabilizing reaction Gibbs free energies are observed for the reactions forming XHgOOH and XHgNOO, where the former is more stable than the latter. The reaction forming IHgNOO has the lowest energy difference among the XHg +Y XHgY reactions.
The following observations emerge concerning the rate of the reactions. The derived rate coefficients (X=Br & X=I cmmolecules) for the reaction X + Hg HgX, were found to be consistent with the literature values, underlining the validity of the computational approach.
@@
The most significant difference between iodine and bromine reaction rates was for reaction R2. The bromine-containing forward reaction rate coefficient (= 8.810 cmmolecules) was calculated to be twice as fast as its iodine counterpart, with a rate of 4.210 cmmolecules. Calculated reaction rates for reaction R2 (XHg +OX XHgOX ) differed in their forward rate coefficients and were between 3.310 and 3.310 cmmolecules. The reaction R2 forming BrHgOBr was about 7.2 times faster than the one forming IHgOI, about 6.5 times faster than the one forming IHgOBr, and about 10 times faster than the one forming BrHgOI.
Atmospheric concentrations of the chemical species of halogen radicals are strongly correlated with local conditions. The general understanding is that there is a larger mixing ratio of bromine radical species in the atmosphere over the Arctic and Antarctic than iodine radical species. Iodine, bromine, and chlorine radicals are implicated in depletion events globally, depending on local conditions.
In accordance of the findings outlined earlier, the reactions between mercury and iodine are not likely to dramatically alter the overall rates of atmospheric reactions involving mercury. This is due to the reactions between mercury and iodine not being faster than those reactions involving bromine, which remains the dominant atmospheric reactant for mercury during springtime in the Arctic and Antarctic. Additionally, the mixing ratio of iodine radical species in the Arctic atmosphere is not greater than that of bromine radical species. These observations reinforce the idea that bromine is the primary atmospheric reactant influencing mercury in the Arctic, Antarctic, and in atmospheric mercury depletion events (ADME). Based on these observations, we can conclude that the AMDEs are dependent on the concentration of these radical species, and the role of iodine species as potential contributors cannot be ruled out. A significant increase in atmospheric concentrations of iodine over the Arctic and Antarctic would only have a substantial impact on mercury’s atmospheric reactions if the atmospheric mixing ratio of iodine markedly surpassed those of atmospheric bromine.

### v2

@@ IV.2 Atmospheric Mixing Ratios in the Arctic and Antarctic
High concentrations of bromine, chlorine, and iodine species have been observed during depletion events in both the Arctic and Antarctic. Maximum mixing ratios of BrO up to 50 ppt in the Antarctic [80] and over 20 ppt at Alert Base in the Arctic [87] have been reported. IO has been measured up to 30 ppt at Mace Head, Ireland [16], and around 20 ppt at Halley Bay in the Antarctic [57]. ClO concentrations reached up to 44 ppt in Barrow, Alaska [19].
Seasonal variations show that BrO mixing ratios at Alert Base ranged from 5 ppt in spring to 1 ppt in summer, rising again in fall [87]. At Scott Base in the Antarctic, BrO increased from below detection limits to about 3 ppt in spring, decreased in summer, and rose slightly in fall. IO levels at Alert Base varied from 0.1 ppt in spring to 0.3 ppt in summer, while at Scott Base, they ranged from 0.35 ppt in spring and fall to 0.15 ppt in summer.
Other radicals like ClO and HO were measured during the SOLVE mission over the Norwegian Sea, showing average ClO concentrations of about 20 ppt in clouds and decreasing from 30 ppt in fall to 10 ppt in winter [67]. HO concentrations dropped from approximately 2 ppt to 0.1 ppt over the same period. Atomic radicals are less frequently reported; Br concentrations ranged from 4 ppt to 14 ppt during ozone depletion events in Barrow, Alaska [81], and I was detected up to 22 ppt at Mace Head [7].
Local conditions greatly influence halogen radical concentrations, affecting atmospheric mercury depletion events (AMDEs). While iodine levels may rise due to global warming, current data indicate that bromine radicals, particularly Br and OBr, are the primary contributors to AMDEs. If iodine species were to become more abundant than bromine, they might compete in atmospheric reactions during mercury depletion events and also play a role in activating bromine [12]. Therefore, although iodine species cannot be dismissed as contributors, the prevalence of bromine radicals suggests they continue to be the dominant factors in AMDEs.
@@ V Conclusion and Outlook
The following observations emerge for the reaction energy. The reactions with the largest stabilization of Gibbs free energy creating Hg(II) from reaction R2 (XHg +Y XHgY or XHg + OY XHgOY) are the reaction forming HgBr2 with a stabilization energy of 61.09 kcal/mol.
The stabilization by the Gibbs free energy decreases with decreasing atomic number, from XHgOI over XHgOBr to XHgOCl. Furthermore, at the other end of the series of molecules, IHgOY molecules are consistently less stable than their BrHgOY counterparts. The least stable XHgOY-product is IHgOCl with a stabilization of 36.06 kcal/mol in Gibbs free energy.
Lower stabilizing reaction Gibbs free energies are observed for the reactions forming XHgOOH and XHgNOO, where the former is more stable than the latter. The reaction forming IHgNO2 has the lowest energy difference among the XHg + OY XHgOY reactions.
The following observations emerge concerning the rate of the reactions.
@@
The most significant difference between iodine and bromine reaction rates was for reaction R2. The bromine-containing forward reaction rate coefficient (= 8.810-13 cm3molecule-1s-1) was calculated to be twice as fast as its iodine counterpart, with a rate of 4.210-13 cm3molecule-1s-1. Calculated reaction rates for reaction R2 (XHg + OY XHgOY ) differed in their forward rate coefficients and were between 3.310-14 and 3.310-15 cm3molecule-1s-1. The reaction R2 forming BrHgOBr was about 7.2 times faster than the one forming IHgOI, about 6.5 times faster than the one forming IHgOBr, and about 10 times faster than the one forming BrHgOI.
In accordance of the findings outlined earlier, the reactions between mercury and iodine are not likely to dramatically alter the overall rates of atmospheric reactions involving mercury. This is due to the reactions between mercury and iodine not being faster than those reactions involving bromine, which remains the dominant atmospheric reactant for mercury. Local conditions can occur where iodine concentrations will be dominant, but in the Arctic and Antarctic, the mixing ratio of iodine radical species in the atmosphere is not greater than that of bromine radical species. These observations reinforce the idea that bromine is the primary atmospheric reactant influencing mercury in the Arctic, Antarctic, and in general ADMEs. Based on these observations, we can conclude that the AMDEs are dependent on the concentration of these radical species, and the role of iodine species as potential contributors cannot be ruled out. A significant increase in local atmospheric concentrations of iodine would only have a substantial impact on mercury’s atmospheric reactions if the atmospheric mixing ratio of iodine markedly surpassed those of atmospheric bromine.